# Axi-majoron for almost everything

### v1

@@ 3.3 Reheating after inflation
$H_{i,\rm eq}=\left(\frac{p_{a,\rm eq}^{i}}{p_{a,*}}\right)^{2}H_{*}=8{\rm B}_{i}^{2}b_{i}H_{*}\ ,$ (3.81)
@@
$\frac{\Gamma_{N_{i}}}{H_{i,\rm eq}}=\frac{3}{4\sqrt{2\lambda_{\phi}}{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\approx\frac{17}{{\rm B}_{i}^{2}\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.82)
If one of RHN mass-eigenstates is responsible for eventual recovering the SM thermal bath, it would have to satisfy $\Gamma_{N_{i}}/H_{i,\rm eq}<0.1\times\left(\Delta N_{\rm eff}^{\rm obs}/0.5\right)^{3/2}$ with $\Delta N_{\rm eff}^{\rm obs}$ being the observational bound on the extra radiation (see Sectoin 5). Combined with the condition in Eq. (3.69), the branching ratio ${\rm B}_{i}$ is then constrained as
${\rm B}_{i}\gtrsim 2.0\times 10^{-2}\left(\frac{0.5}{\Delta N_{\rm eff}^{\rm obs}}\right)^{4/5}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)^{8/15}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{2/5}\ .$ (3.83)
Hence, except the case of $\phi\sim 10^{9}{\rm GeV}$ with ${\rm B}_{i}=\mathcal{O}(0.1)$, typically heavy states are expected to decay around or before the epoch they become non-relativistic, and never dominate the universe. The standard thermal bath in this case can be recovered by the decay of the lightest RHN mass-eigenstate $N_{1}$ if $m_{\nu_{1}}$(or equivalently $y_{\nu_{1}}$) is sufficiently smaller than $m_{\nu}$.
Note that, as inferred from Eq. (3.83), if $\Delta N_{\rm eff}^{\rm obs}=\mathcal{O}(0.1)$ with $\phi_{0}\ll\phi_{0}^{\rm ref}$, heavy RHN states can decay while they dominate the universe, and recover the standard thermal bath by themselves. The lightest RHN $N_{1}$ in such a case can become the main component of dark matter, having self-interactions mediated by a very light axi-majoron. Although $N_{1}$ should have essentially zero Yukawa coupling $y_{\nu_{1}}$, in principle it is a possibility one can consider. We do not discuss this case further in this work, though.
@@
$f_{N}(\xi_{\phi},\phi_{0})\equiv\left[\frac{H_{*}}{\Gamma_{N}/2}\left(\frac{m_{N}}{m_{\phi}}\right)^{2}\right]^{1/3}=\left[\frac{2\sqrt{2}\lambda_{\phi}^{1/2}}{3}\left(\frac{v_{h}^{2}}{m_{\nu_{N}}\phi_{0}}\right)\right]^{1/3}\simeq 0.3\xi_{\phi}^{1/3}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\ ,$ (3.85)
@@
where we used Eq. (3.57) with $g_{*}(T_{\gamma,N})=106.75$ in the second line. Hence, unless ${\rm B}_{N}\ll 10^{-4}$, the heavy states $N_{2,3}$ can be thermalized with the SM thermal bath via inverse decay processes. Note that $m_{\phi}/T_{\gamma,N}\simeq 6.7\times f_{N}^{-1/2}\left(10^{-2}/b_{N}\right)^{3/4}$ with $b_{N}\equiv b_{3}\approx b_{2}$. Hence, if $b_{N}=\mathcal{O}(0.01-0.1)$, the inverse decay, $N+N\to\phi$, can also be efficient enough to thermalize $\phi$-particles. But the impact of the process is minor in our discussion. Axi-majorons are not thermalized even in this case, since the interaction is suppressed at least by $\mathcal{O}(T_{\gamma,N}^{2}/\phi_{0}^{2})$. There would be thermal contribution to the energy of axi-majorons from partial decays of thermalized $\phi$-particles, but it is subdominant.
@@ 4 Leptogenesis
Thanks to the symmetry-breaking $\xi_{a}$-term in Eq. (2.7), the inflation along $\Phi$-direction can generate angular motion at the end of inflation, as long as $\theta_{\rm ini}$ is not very close to zero and the mass scale along the angular direction is somewhat smaller but not extremely smaller than the expansion rate during inflation. Since $U(1)_{\rm PQ}$ is conserved in ${\tilde{V}}$, the angular momentum of $\Phi$ at the end of inflation can be conserved for $\phi\ll M_{\rm P}$, and corresponds to the Peccei-Quinn number asymmetry stored in $\Phi$:
@@
$T_{e}\approx\left(\frac{\pi^{2}}{90}g_{*}^{\Phi}\right)^{-1/4}\sqrt{H_{e}M_{\rm P}}=\left(\frac{2\pi^{2}g_{*}^{\Phi}}{15}\right)^{-1/4}\lambda_{\phi}^{1/4}\phi_{e}\ ,$ (4.4)
@@
It seems that $Y_{\Phi,e}$ depends only on $\alpha$, when $\lambda_{\phi}$ and $\xi_{\phi}$ satisfy the relation in Eq. (3.51) to provide inflationary observables matching observations. However, it should be noted that Eq. (4.5) provides just a qualitative idea of the maximal possibility of $Y_{\Phi,e}$. The exact parametric dependence of $Y_{\Phi,e}$ is non-trivial, since it involves the motion of $\Phi$ at the end of inflation that is affected by the initial position of $\phi$, i.e., $\phi_{\rm ini}$. For later convenience, we may express $\dot{\theta}_{e}$ in terms of $Y_{\Phi,e}$ as
@@
The eventual decays of $\Phi$ field transfer the PQ-asymmetry to the asymmetry of RHN-helicity states via the interaction $y_{N}$ responsible for generating the majorana mass of RHNs. One may assign 2-units of lepton number to the Peccei-Quinn field, and one-unit to the right-handed neutrino $\nu_{R}$. Then, both interactions $y_{N}$ and $y_{\nu}$ respects not only $U(1)_{\rm PQ}$ but also $U(1)_{B-L}$, and the PQ-number asymmetry of $\Phi$ field can be regarded as the lepton number asymmetry as $\Delta n_{L}^{\Phi}=2\Delta n_{\Phi}$. In the broken phase of $U(1)_{\rm PQ}$-symmetry, the non-zero majorana mass-term breaks the lepton number conservation. However, as eigenstates of the equation of motion, helicity states are not affected by the majorana mass-term and the asymmetry between the helicity states of RHNs is not changed. The Yukawa interaction $y_{\nu}$ violates the helicity conservation, but the effect depends on the speed of RHNs. The branching ratio of the channel, $N_{i}^{+}\to\ell+H$, is $\left(1+\beta_{i}\right)/2$ whereas that of the channel, $N_{i}^{+}\to\overline{\ell}+\overline{H}$, is $\left(1-\beta_{i}\right)/2$ with $\beta_{i}$ being the speed of $N_{i}$. Hence, the helicity asymmetry of $N_{i}$ induced from the decay of $\Phi$-field is only partially transferred to the visible sector by the suppression factor $\beta_{i}$. As a result, if $N_{i}$ is never equilibrated, its decay contributes to the lepton number asymmetry of the visible sector as [60]
$\Delta n_{L}^{i}=\Delta n_{\rm hel}^{i}\times\beta_{i}$ (4.7)
where $\Delta n_{\rm hel}^{i}$ denotes the asymmetry between $N_{i}^{+}$ and $N_{i}^{-}$.
The late time asymmetry in the visible sector depends not only on the speeds of RHNs but also on the background temperature of the SM thermal bath when RHNs decay. Depending on parameters causing different situations, there are several possibilities for generating baryon number asymmetry of the universe: Affleck-Dine (AD) baryogenesis [27] through the decay chain such as $\Phi\to\nu_{R}\to\ell$ with RHNs out of equilibrium, spontaneous leptogenesis thanks to the motion of the axi-majoron field with thermal RHNs [61], and resonant leptogenesis [28] from decays of non-thermal RHNs.
@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
A part of the PQ-number asymmetry can be transferred to the visible sector through the helicity states of $N_{1}$, while the contributions from heavier states are washed to a much smaller value. It is possible for the heavy states to decay when they just become non-relativistic while producing SM particles with a temperature $T\sim m_{N}\gtrsim m_{N_{1}}$. For a diagonal mass matrix of RHNs, which we are assuming in this work, the scattering and inverse process to thermalize $N_{1}$ is efficient only when $T\sim m_{N_{1}}$ with $T$ being the temperature in the standard radiation-dominated universe. In such a thermal bath, the scattering rate of $N_{1}$ to SM particles when $T\gtrsim m_{N_{1}}$ is expected to be $\Gamma_{1,s}=c_{1,s}y_{\nu_{1}}^{2}T$ with $c_{1,s}=\mathcal{O}(10^{-2})$ [60] being a numerical factor including the effect of SM couplings and the number of allowed channels. Note that the expansion rate at $T=m_{N_{1}}$ is given by
$H(T=m_{N_{1}})=H_{\rm SM}(T=m_{N_{1}})/\sqrt{2{\rm B}_{N}}$
with $H_{\rm SM}(T)$ being the expansion rate in the standard cosmology in a radiation-dominated universe with temperature $T$. The ratio of $\Gamma_{1,s}$ to the expansion rate is then found as
$\frac{\Gamma_{1,s}}{H(T=m_{N_{1}})}$ $=$ $2\left(\frac{90}{\pi^{2}g_{*}(T=m_{N_{1}})}\right)^{1/2}c_{1,s}\sqrt{2{\rm B}_{N}}\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}$ (4.8)
$\approx$ $5.2\times\left(\frac{c_{1,s}}{10^{-2}}\right)\left(\frac{{\rm B}_{N}}{0.1}\right)^{1/2}\left(\frac{m_{\nu_{1}}}{m_{\nu}}\right)\ ,$
where $g_{*}(T=m_{N_{1}})=106.75$ was used in the second line. Hence, for $m_{\nu_{1}}$ smaller than $m_{\nu}$ by a couple of orders of magnitude that may be necessary in any case for $N_{1}$ to recover the SM thermal bath finally (see Eq. (3.82)), the scattering of $N_{1}$ can be inefficient. Then, even though there would be thermal contribution to the abundance of $N_{1}^{\pm}$s as a subdominant portion, once produced in the decay of the inflaton, $N_{1}$ can be never thermalized with the SM particles.
@@
$\left(\Delta n_{\rm hel}^{1}\right)_{\times}\approx\frac{4c_{\rm AD}{\rm B}_{1}}{5}\frac{\Gamma^{\phi}}{H_{\times}}\left(\Delta n_{\Phi}\right)_{\times}\ .$ (4.9)
@@
$\Delta_{\rm AD}\equiv\left(c_{\rm AD}\frac{2\Gamma^{\phi}}{5H_{\times}}\right)\ ,$ (4.10)
@@
relative to the expectation in conventional scenarios. For the decoupled $N_{1}$, the speed $\beta_{1}$ at its decay is given by
$\beta_{1}=\left(\frac{H_{1,\rm eq}}{H_{1,\rm non}}\right)^{1/2}\left(\frac{H_{\rm d}}{H_{1,\rm eq}}\right)^{2/3}=\left({\rm B}_{1}\lambda_{\phi}\right)^{-1/3}\left(\frac{m_{\nu_{1}}\phi_{0}}{v_{h}^{2}}\right)^{2/3}\ .$ (4.11)
@@
Then, the late time asymmetry transferred through the decays of $N_{1}$s is found to be
$Y_{B,{\rm AD}}^{<}$ $=$ $-\Delta_{\rm AD}\times\beta_{1}\times\frac{12}{37}\left(\frac{\sqrt{3}}{8\pi}\right)^{1/2}\left(\frac{g_{*}^{\Phi}}{g_{*}(T_{\rm d})}\right)^{1/4}\lambda_{\phi}^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}Y_{\Phi,e}$
$\simeq$ $-1.7\times 10^{-10}\times\left(\frac{c_{\rm AD}}{0.1}\right)\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{\xi_{\phi}}{10}\right)^{11/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\left(\frac{Y_{\Phi,e}}{10^{-1}}\right)\ ,$

### v2

@@ 3.3 Reheating after inflation
${\tilde{H}}_{i,\rm eq}=\sqrt{2}\left(\frac{E_{a,\rm eq}^{i}}{{\tilde{E}}_{a,*}}\right)^{2}{\tilde{H}}_{*}=2\sqrt{2}{\rm B}_{i}^{2}b_{i}{\tilde{H}}_{*}\ ,$ (3.82)
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm eq}}=\frac{3}{2\sqrt{\lambda_{\phi}}{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\approx\frac{51}{{\rm B}_{i}^{2}\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.83)
If one of RHN mass-eigenstates is responsible for eventual recovering the SM thermal bath, it would have to satisfy $\Gamma_{N_{i}}/{\tilde{H}}_{i,\rm eq}<0.1\times\left(\Delta N_{\rm eff}^{\rm obs}/0.5\right)^{3/2}$ with $\Delta N_{\rm eff}^{\rm obs}$ being the observational bound on the extra radiation (see Sectoin 5). Combined with the condition in Eq. (3.71), the branching ratio ${\rm B}_{i}$ is then constrained as
${\rm B}_{i}\gtrsim 0.40\times\left(\frac{0.5}{\Delta N_{\rm eff}^{\rm obs}}\right)^{3/4}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/4}\ ,$ (3.84)
but it can not be larger then $1/3$ as shown in Eq. (3.61). Hence, the standard thermal bath in this case is to be recovered by the decay of the lightest RHN mass-eigenstate $N_{1}$ with $m_{\nu_{1}}$ satisfying the following condition:
$\frac{m_{\nu_{1}}}{m_{\nu}}<2\times 10^{-4}\times\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{\xi_{\phi}}{10}\right)\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.85)
Then, it is necessary to have $\phi_{0}\sim\phi_{0}^{\rm ref}$ for the observed relic density of dark matter (see Section 5), and heavy states are expected to decay around or before the epoch they become non-relativistic, and never dominate the universe.
@@
$f_{N}(\xi_{\phi},\phi_{0})\equiv\left[\frac{{\tilde{H}}_{*}}{\Gamma_{N}/2}\left(\frac{m_{N}}{m_{\phi,0}}\right)^{2}\right]^{1/3}\simeq 0.3\xi_{\phi}^{1/3}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\ ,$ (3.87)
@@
where we used Eq. (3.57) with $g_{*}(T_{\gamma,N})=106.75$ in the second line. Hence, unless ${\rm B}_{N}\ll 10^{-4}$, the heavy states $N_{2,3}$ can be thermalized with the SM thermal bath via inverse decay processes. Note that $m_{\phi,0}/T_{\gamma,N}\simeq 6.7\times f_{N}^{-1/2}\left(10^{-2}/b_{N}\right)^{3/4}$ with $b_{N}\equiv b_{3}\approx b_{2}$. If $b_{N}=\mathcal{O}(0.01-0.1)$, the inverse decay, $N+N\to\phi$, can also be efficient enough to thermalize $\phi$-particles. But the impact of the process is minor in our discussion. Axi-majorons are not thermalized even in this case, since the interaction is suppressed at least by $\mathcal{O}(T_{\gamma,N}^{2}/\phi_{0}^{2})$. There would be thermal contribution to the energy of axi-majorons from partial decays of thermalized $\phi$-particles, but it is subdominant.
Case of $\phi_{*}>\phi_{\times}$:
If $\phi_{*}>\phi_{\times}$, i.e., $\xi_{\phi}>100\sqrt{\phi_{0}/\phi_{0}^{\rm ref}}$ (see Eq. (3.63)), $\Phi$ field decays before falling into the region of the true vacuum position. Even in this case, $N_{i}$ can be kept out of kinetic or thermal equilibrium as long as $\xi_{\phi}\sim\mathcal{O}(10^{2})$ with $\phi_{0}\approx\phi_{0}^{\rm ref}$. The expansion rate at the epoch when $N_{i}$ becomes non-relativistic is found as
$H_{i,\rm non}=\left(\frac{2m_{N_{i}}}{m_{\phi,*}}\right)^{2}H_{*}$ (3.89)
where $m_{\phi,*}=\sqrt{3\lambda_{\phi}}\phi_{*}$ and $H_{*}=\lambda_{\phi}^{1/2}\phi_{*}^{2}/2\sqrt{3}M_{\rm P}$ with $\phi_{*}$ given by Eq. (3.63). The energy density of $N_{i}$ starts dominating when $H=H_{i,\rm eq}$ with
$H_{i,\rm eq}=\sqrt{2}{\rm B}_{i}^{2}H_{i,\rm non}=\sqrt{2}{\rm B}_{i}^{2}\left(\frac{2m_{N_{i}}}{m_{\phi,*}}\right)^{2}H_{*}\ .$ (3.90)
Note that $H_{1,\rm eq}$ is always smaller than ${\tilde{H}}_{1,\rm eq}$. Then, one finds
$\frac{\Gamma_{N_{i}}}{H_{i,\rm eq}}=\frac{3\sqrt{3\lambda_{\phi}/2}}{8\pi{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}M_{\rm P}}{v_{h}^{2}}\right)>0.25\times\left(\frac{1/3}{{\rm B}_{i}}\right)^{2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/2}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)$ (3.91)
where the last inequality is from the condition, $\phi_{*}>\phi_{\times}$. Hence, the heavy RHN states can not be responsible for the late time thermal background consisting of the SM particles in this case, too.
The constraint on $m_{\nu_{1}}$ to recover the SM thermal bath without too much dark radiation in this case is found as
$\frac{m_{\nu_{1}}}{m_{\nu}}\lesssim 10^{-3}\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{100}{\xi_{\phi}}\right)\ .$ (3.92)
Therefore, irrespective of whether $\phi_{*}$ is smaller than $\phi_{\times}$ or not, $m_{\nu_{1}}$ should be smaller than $m_{\nu}$ by several orders of magnitude.
@@ 4 Leptogenesis
Thanks to the symmetry-breaking $\xi_{a}$-term in Eq. (2.7), the inflation along $\Phi$-direction can generate angular motion at the end of inflation, as long as $\theta_{\rm ini}$ is not very close to zero and the mass scale along the angular direction is somewhat smaller but not extremely smaller than the expansion rate during inflation  [^19] This is a kind of inflationary Affleck-Dine mechanism, and various different realizations have been considered (see for example Refs. [63, 64, 65, 66, 67]).. Since $U(1)_{\rm PQ}$ is conserved in ${\tilde{V}}$, the angular momentum of $\Phi$ at the end of inflation can be conserved for $\phi\ll M_{\rm P}$, and corresponds to the Peccei-Quinn number asymmetry stored in $\Phi$:
@@
$T_{e}\equiv\left(\frac{\pi^{2}}{90}g_{*}^{\Phi}\right)^{-1/4}\sqrt{H_{e}M_{\rm P}}=\left(\frac{2\pi^{2}g_{*}^{\Phi}}{15}\right)^{-1/4}\lambda_{\phi}^{1/4}\phi_{e}\ ,$ (4.4)
@@
It seems that, when $\lambda_{\phi}$ and $\xi_{\phi}$ satisfy the relation in Eq. (3.51) to provide inflationary observables matching observations, $Y_{\Phi,e}$ depends only on $\alpha$, modulo the dependence on the initial condition, $\theta_{\rm ini}$. It should be noted that Eq. (4.5) provides just a qualitative idea of the maximal possibility of $Y_{\Phi,e}$, though. The exact parametric dependence of $Y_{\Phi,e}$ is non-trivial, since it involves the motion of $\Phi$ during inflation that is affected by the initial position of $\phi$, i.e., $\phi_{\rm ini}$ in addition to $\theta_{\rm ini}$. For later convenience, we may express $\dot{\theta}_{e}$ in terms of $Y_{\Phi,e}$ as
@@
The eventual decays of $\Phi$ field transfer the PQ-asymmetry to the asymmetry of RHN-helicity states via the interaction $y_{N}$ responsible for generating the majorana mass of RHNs. One may assign 2-units of lepton number to the Peccei-Quinn field, and one-unit to the right-handed neutrino $\nu_{R}$. Then, both interactions $y_{N}$ and $y_{\nu}$ respects not only $U(1)_{\rm PQ}$ but also $U(1)_{B-L}$, and the PQ-number asymmetry of $\Phi$ field can be regarded as the lepton number asymmetry as $\Delta n_{L}^{\Phi}=2\Delta n_{\Phi}$. In the broken phase of $U(1)_{\rm PQ}$-symmetry, the non-zero majorana mass-term breaks the lepton number conservation. However, as eigenstates of the equation of motion, helicity states are not affected by the majorana mass-term and the asymmetry between the helicity states of RHNs is not changed, especially, in the CMB frame of the homogeneous and isotropic universe [68]. The Yukawa interaction $y_{\nu}$ violates the helicity conservation, but the effect depends on the speed of RHNs. The branching ratio of the channel, $N_{i}^{+}\to\ell+H$, is $\left(1+\beta_{i}\right)/2$ whereas that of the channel, $N_{i}^{+}\to\overline{\ell}+\overline{H}$, is $\left(1-\beta_{i}\right)/2$ with $\beta_{i}$ being the speed of $N_{i}$. Hence, the helicity asymmetry of $N_{i}$ induced from the decay of $\Phi$-field is only partially transferred to the visible sector by the suppression factor $\beta_{i}$. As a result, if $N_{i}$ is never equilibrated, its decay contributes to the lepton number asymmetry of the visible sector as [68]
$\Delta n_{L,i}=\Delta n_{i}\times\beta_{i}$ (4.7)
where $\Delta n_{i}$ denotes the helicity asymmetry between $N_{i}^{+}$ and $N_{i}^{-}$.
The late time asymmetry in the visible sector depends not only on the speeds of RHNs but also on the background temperature of the SM thermal bath when RHNs decay. Depending on parameters causing different situations, there are several possibilities for generating baryon number asymmetry of the universe: Affleck-Dine (AD) baryogenesis [29] through the decay chain such as $\Phi\to\nu_{R}\to\ell$ with RHNs out of equilibrium, spontaneous leptogenesis thanks to the background motion of the axi-majoron field with thermal RHNs [69], and resonant leptogenesis [30] from decays of non-thermal RHNs.
@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
A part of the PQ-number asymmetry can be transferred to the visible sector through the helicity states of $N_{1}$, while the contributions from heavier states are washed to a much smaller value. It is possible for the heavy states to decay when they just become non-relativistic while producing SM particles with a temperature $T\gtrsim m_{N}\gtrsim m_{N_{1}}$. In such a thermal bath, the scattering rate of $N_{1}$ to SM particles when $T\gtrsim m_{N_{1}}$ is expected to be $\Gamma_{1,s}=c_{1,s}y_{\nu_{1}}^{2}T$ with $c_{1,s}=\mathcal{O}(10^{-2})$ [68] being a numerical factor including the effect of SM couplings and the number of allowed channels. The expansion rate at $T=m_{N_{1}}$ in our scenario is given by
$H(T=m_{N_{1}})=H_{\rm SM}(T=m_{N_{1}})/\sqrt{2{\rm B}_{N}}$ (4.8)
with $H_{\rm SM}(T)$ being the expansion rate in the standard cosmology in a radiation-dominated universe with temperature $T$. It gives
$\frac{\Gamma_{1,s}}{H(T=m_{N_{1}})}$$\approx$ $5.2\times\left(\frac{c_{1,s}}{10^{-2}}\right)\left(\frac{{\rm B}_{N}}{0.1}\right)^{1/2}\left(\frac{m_{\nu_{1}}}{m_{\nu}}\right)\ ,$ (4.9)
where $g_{*}(T=m_{N_{1}})=106.75$ was used for the numerical approximation. Meanwhile, the freeze-out temperature($T_{1,\rm fo}$) of the dominant inverse processes washing-out the $N_{1}$’s contribution to the visible sector lepton number asymmetry can be somewhat lower than $m_{N_{1}}$. It can be found by equating the expansion rate to the rate of the processes ($\gamma_{1,\rm ID}$) [69]:
$H(T_{1,\rm fo})=\gamma_{1,\rm ID}\equiv\frac{n_{N_{1}}^{({\rm eq})}}{n_{\ell}^{({\rm eq})}}\frac{K_{1}(x_{1,\rm fo})}{K_{2}(x_{1,\rm fo})}\Gamma_{N_{1}}$ (4.10)
where $x_{1,\rm fo}\equiv m_{N_{1}}/T_{1,\rm fo}$, $n^{({\rm eq})}$ is the equilibrium number density, and $K_{1,2}$ are the modified Bessel function. It turns out that $x_{1,\rm fo}$ depends only on $m_{\nu_{1}}$, and the dependence is shown in Fig. 9.
Figure 9: The freeze-out temperature of thermal $N_{1}$ in terms of $x_{1,\rm fo}$ as a function of $m_{\nu_{1}}$.
From the figure, one sees that $N_{1}$ would be never in equilibrium if $m_{\nu_{1}}/m_{\nu}<6\times 10^{-2}$. Note that the upper limit is already well above the bound in Eq. (3.85) for the late time recovering of the SM thermal bath without too much dark raditaion. Hence, as long as Eq. (3.85) is satisfied, even though there would be thermal contribution to the abundance of $N_{1}^{\pm}$s as a subdominant portion, $N_{1}$ produced in the decay of the inflaton can be never thermalized with the SM particles.
@@
$\Delta n_{1,\times}\approx\frac{4c_{\rm AD}{\rm B}_{1}}{5}\frac{\Gamma^{\phi}}{H_{\times}}\Delta n_{\Phi,\times}\ .$ (4.11)
@@
$\Delta_{\rm AD}\equiv c_{\rm AD}\times\frac{2\Gamma^{\phi}}{5H_{\times}}\ ,$ (4.12)
@@
relative to the expectation in conventional scenarios. For the decoupled $N_{1}$, its speed $\beta_{1,\rm d}^{<}$ when it decays is given by
$\beta_{1,\rm d}^{<}$ $=$ $\frac{m_{\phi,\times}}{2m_{N_{1}}}\left(\frac{{\tilde{H}}_{*}}{H_{\times}}\right)^{2/3}\left(\frac{{\tilde{H}}_{1,\rm eq}}{{\tilde{H}}_{*}}\right)^{1/2}\left(\frac{H_{\rm d}}{{\tilde{H}}_{1,\rm eq}}\right)^{2/3}$ (4.13)
$\simeq$ $10^{-3}\times\left(\frac{\xi_{\phi}}{10}\right)^{2/3}\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{2/3}\ ,$
where $m_{\phi,\times}\equiv\sqrt{3\lambda_{\phi}}\phi_{\times}$. Also, the yield of the helicity asymmetry of $N_{1}$ at the epoch is found as (see Appendix E)
$Y_{1,\rm d}^{<}=4\Delta_{\rm AD}\left(\frac{{\tilde{H}}_{*}}{{\tilde{H}}_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,\times}}Y_{\Phi,e}\ .$ (4.14)
@@
Then, the late time asymmetry transferred through the decays of $N_{1}$s is found to be [70]
$Y_{B,{\rm AD}}^{<}$ $=$ $\frac{12}{37}\times\beta_{1,\rm d}^{<}Y_{1,\rm d}^{<}$ (4.15)
$\simeq$ $8.4\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi}}{10}\right)^{19/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\left(\frac{Y_{\Phi,e}}{10^{-2}}\right)\ ,$